OpenStudy (anonymous):

Given f(x)=4-x^2 for x is less than or equal to 0, find f^-1(x).

OpenStudy (anonymous):

first replace f(x) with y y=4-x^2 now replace each x with y and each y with x x=4-y^2 now solve for y

OpenStudy (anonymous):

so y = sqrt(4-x) ?

OpenStudy (anonymous):

yes replace y with f^-1(x) !!

OpenStudy (anonymous):

Ok thank you!